OpenStudy (anonymous):

in a class,the ratio of the number of boys to the number of girls is 3:5 .If there are 18 boys , how many girls are there?

OpenStudy (katrinakaif):

Set a proportion: 3 18 - = -- 5 x Cross Multiply: 3x = 18(5) and Solve.

OpenStudy (anonymous):

3/5=18/x x=(18*5)/3 x=30

OpenStudy (anonymous):

Katrina kaif the answer is nt good :)

OpenStudy (wasiqss):

hehe

OpenStudy (katrinakaif):

I am not here to give you the answer.

OpenStudy (anonymous):

so why did u answer my question ...............

OpenStudy (katrinakaif):

I gave you a solution.

OpenStudy (anonymous):

3/5=18/x this implies x=30

OpenStudy (anonymous):

and katrina kaif i have talk to u well